# Smoothness and other hyperparameter estimation for inverse problems related to data assimilation

### v1

@@ 3.2 Metropolis-within-Gibbs
One could use either direct sampling from each conditional above or if this is not possible a small number of Metropolis iterations targeting each of the above densities. [^1] One may even go further and implement MwG for each coordinate or blocks of coordinates of the variables in $v,\theta,\phi)$. The latter is referred to as MwG. Sampling directly using conjugate relationships is recommended when available to accelerate sampling.
@@ 4.1 Case study 1: inverse problem for Navier-Stokes
We are interested in comparing pCN and MwG for similar experiments to [26], where $\alpha$ is assumed to be known and set to either $2$ or $2.2$. We produce figures in the study to compare the performance in initial condition estimation of the MwG to pCN, assuming correct and incorrect values of $\alpha$. For the PDE solver, we will use a spectral Galerkin method. We consider the finite dimensional projection onto the Hilbert space spanned by $\{\psi_{k}\}_{k\in\mathbb{L}_{n}}$ where $n\in\mathbb{N}\setminus\{0\}$ and $\mathbb{L}_{n}=\{(k_{1},k_{2})\in\mathbb{Z}^{2}\setminus\{0\}:\text{max}(|k_{1}|,|k_{2}|)<n/2\}$. We refer to the size of the implied mesh, $d_{u}$, as the dimensionality of u. We observe that $A$ is diagonalized in $U$ with eigenvalues $\{\lambda_{k}\}_{k\in\mathbb{Z}^{2}\setminus{\{0\}}}$, where $\lambda_{k}=|k|^{2}$. Periodic boundary conditions allow the use of spectral Galerkin methods when solving the Navier-Stokes PDE numerically with Fast Fourier Transforms (FFTs). When facing different boundary conditions or forcing, we would use Finite Elements techniques or other methods [21]. Convolutions arising from products in the nonlinear term are computed via FFTs on our mesh with additional antialiasing using double-sized zero padding. In addition, exponential time differencing is used as in [12] to deal with the resulting stiff system, whereby an analytical integration is used for the linear part, together with an explicit numerical Euler integration scheme for the nonlinear part.
@@
Figure 2: (Stationary regime) Vorticity of the initial condition $v_{0}$ with the truth (far left) and MCMC estimates: knowing the true $\alpha=2.2$ value (centre left), underestimating with $\alpha=1.5$ (centre), overestimating with $\alpha=3$ (centre right) and estimating $\alpha$ using MwG (far right). The crosses indicate the positions where the vector field is observed. We observe similar estimates from the algorithms knowing the correct smoothness and estimating it, albeit giving a smoother estimate.
The diagnostics plots for pCN and MwG are in the Appendix in section C.1. Run times are shown in Table C. We remark that as expected the runtimes are close to that of Algorithm 1, with $\alpha,\beta^{2}$ known. In Fig. 1 we plot the trace plots and densities of our parameter estimation and the results prove the success of the parameter estimation even with uninformative priors. We now study the algorithm’s uncertainty quantification of the initial condition. In Fig. 2 we plot the vorticity of the initial condition’s estimates against the true value. We observe similar estimates from the algorithms knowing the correct $\alpha$ and estimating it. The estimates recover the true initial condition relatively accurately, albeit giving a smoother estimate.
Figure 3: (Stationary regime) Heat map of the variance of the vorticity of $v_{0}$ vs $\alpha$ (top) in the stationary setting, as previously arranged, with: $\alpha=2.2$ value (far left), $\alpha=1.5$ (centre left), $\alpha=3$ (centre right) and estimating $\alpha$ using MwG (far right). Note the scale of the variance of $v_{0}$’s estimates being inversely proportional to $\alpha$. We also plot the trajectory of the velocity in the 2nd dimension of those estimates at observed (middle) and unobserved (bottom) locations. The vertical dashed line indicates the time of the last observation. The algorithms knowing the true $\alpha$ and estimating it via MwG exhibit similar behaviour and uncertainty quantification.
We now plot in Fig. 3 the variance of the estimates’ vorticity to fully evaluate the algorithm’s uncertainty quantification. We see from the scales that the variance is highly impacted by the choice of $\alpha$. We recover the order of variance when the true $\alpha$ is known only with the MwG. When $\alpha$ is over and underestimated the uncertainty is under and over estimated, respectfully. Practitioners often choose to sacrifice uncertainty quantification for improved point estimates and vice-versa. This could be used within our framework by adjusting the priors of $\alpha$. Note however we will use a high viscosity and this lowers the system’s energy, and in doing so, this stationary regime renders our variance estimation less crucial and the predictive power of all the algorithms similar after a few timesteps. We repeat this below for the ”chaotic” regime with lower viscosity and more frequent and informative observations to investigate the differences.
@@
Figure 4: (Chaotic regime) Heat map of the variance of the velocity squared of $v_{0}$ vs $\alpha$ (top) in the chaotic setting, as previously arranged, with: $\alpha=2.2$ value (far left), $\alpha=1.5$ (centre left), $\alpha=3$ (centre right) and estimating $\alpha$ using MwG (far right). The white circles in the top panels indicate the positions where the vector field is observed. We also plot the trajectory of the velocity in the 2nd dimension of those estimates at observed (middle) and unobserved (bottom) locations. The vertical dashed line indicates the time of the last observation. Note the scale of the variance of the prediction estimates being inversely proportional to $\alpha$.
@@
We obtain similarly good results in parameter estimation as in the stationary case, see Fig. 9. This is also the case for the vorticity point estimates shown in Fig. 10. These plots can be found in the Appendix in section C.2. In Fig. 4 we present the results as in Fig. 3 previously. The scale of the variance of the vorticity estimates is as before at unobserved locations, however there is spatial correlation introduced by the observations. The likelihood is peaky around the informative observation locations, which reduces the variance of the posterior of the velocity at these locations, as well as the posterior’s dependence on the prior. For higher values of $\alpha$ this spatial correlation is more pronounced as the corresponding prior variance is smoothed. So a higher $\alpha$ leads to a less informative prior variance, additionally increasing the posterior’s dependence on the likelihood. This is why the variance is impacted by $\alpha$ as it was in the stationary setting only at unobserved locations. This can be additionally seen in the plots of the trajectories of the velocity’s estimates at observed times. Due to the chaotic regime having high energy, the trajectory plots show that the difference in variances does not vanish in time and in turn impacts the estimate’s predictive power. $\alpha=3$ leads to an underestimated variance at unobserved locations: the confidence interval does not capture the true value after very few time steps past the observations. At observed locations the forecast performs well, meaning the likelihood is enough to correct the misspecified prior. As for $\alpha=1.5$, the variance is overestimated leading to poor uncertainty quantification at all times. At observed locations its forecast performs well until the last observation, its predictive power degrades quickly as the variance explodes. Then pCN with $\alpha=1.5$ performs poorly at uncertainty quantification, particularly for unobserved states and prediction. This shows that even where the likelihood is most influential, the prior still influences the algorithm’s predictive power. Finally the MwG reproduces similar performance to the algorithm assuming the correct value of $\alpha$ and performs similarly well even in prediction.
@@ C.1 Diagnostics
Figure 8: MwG diagnostics: the cumulative average of the acceptance ratio of the initial condition and of $\alpha$ is around $0.23$ as desired. The acceptance ratio is diverse as desired. The trace plot of the real part of the velocity in the first dimension at various points shows each MCMC estimate (blue) captures the true value (green) in its distribution, with its cumulative average (red) converging to a value mostly close to the truth.
Fig. 7 diagnostics for pCN show that $N=8\cdot 10^{5}$ iterations are sufficient for the MCMC to provide good estimates. In Fig. 7(a) we can observe that the acceptance ratio is stable after the burn-in period of $8\cdot 10^{4}$ and between the desired values of 0.2 and 0.3. Focusing on the real value of a single Fourier coefficient, in Fig. 7(b) the MCMC estimate converges to 0.54-0.54 very close to the true value 0.56-0.56. Additionally its autocorrelation plot in Fig. 7(c) suggests that despite high decorrelation lags, the high number of MCMC iterations will lead to a reasonable effective sample size. For MwG, we increase the number of MCMC moves to the traditional MCMC algorithm to take into account the Gibbs sampler’s moves on the hyperparameters: $N=2.4\cdot 10^{6}$ for $p_{v}=p_{\beta^{2}}=p_{\alpha}=\frac{1}{3}$. To fully observe the behaviour of the parameter estimation, we use longer runs: $N=6\cdot 10^{6}$. This is solely for our study and would not be needed in practice. We found the optimal step-size for $\alpha$ to be $0.96$ for both regimes. Trace plots of the velocity estimates are shown in Fig. 8.
@@
MwG Stationary N-S $N=6\cdot 10^{6}$ 85 hours $2\cdot 10^{8}$
@@
MwG Chaotic N-S $N=6\cdot 10^{6}$ 49 hours $5\cdot 10^{6}$
@@
Table 1: Run times of the MwG algorithm estimating $\theta,\phi$ and of the pCN algorithm knowing $\theta,\phi$. The MwG with $p_{u}=p_{\beta^{2}}=p_{\alpha}=1/3$ is slightly slower than pCN due to additional memory requirements for storing hyperparameter samples. Taking this into account and looking at the number of times the PDE solver was ran, the run times highlight that estimating $\theta$ does not significantly increase the run times over the known $\theta$ case. The computational run times are dominated by the runs of the PDE solver when evaluating the likelihood, which was confirmed when profiling the code.
C.2 Chaotic regime plots
Figure 10: Vorticity of the initial condition $v_{0}$ with the truth (far left) and MCMC estimates: knowing the true $\alpha=2.2$ value (centre left), underestimating $\alpha=1.5$ (centre), overestimating $\alpha=3$ (centre right) and estimating $\alpha$ using MwG (far right). The crosses indicate the positions where the vector field is observed. We observe similar estimates from the algorithms knowing the correct smoothness and estimating it, albeit giving a smooth estimate.

### v2

@@ 3.2 Metropolis-within-Gibbs
One could use either direct sampling from each conditional above or if this is not possible a small number of Metropolis iterations targeting each of the above densities. [^1] One may even go further and implement MwG for each coordinate or blocks of coordinates of the variables in $(v,\theta,\phi)$. The latter is referred to as MwG. Sampling directly using conjugate relationships is recommended when available to accelerate sampling.
@@ 4.1 Case study 1: inverse problem for Navier-Stokes
We are interested in comparing MwG where $\alpha$ is fixed and estimated for similar experiments to [26], where $\alpha$ is assumed to be known and set to either $2$ or $2.2$. We produce figures in the study to compare the performance in initial condition $v_{0}$ and parameter $\beta^{2}$ estimation of the MwG assuming correct and incorrect values of $\alpha$ and estimating $\alpha$. For the PDE solver, we will use a spectral Galerkin method. We consider the finite dimensional projection onto the Hilbert space spanned by $\{\psi_{k}\}_{k\in\mathbb{L}_{n}}$ where $n\in\mathbb{N}\setminus\{0\}$ and $\mathbb{L}_{n}=\{(k_{1},k_{2})\in\mathbb{Z}^{2}\setminus\{0\}:\text{max}(|k_{1}|,|k_{2}|)<n/2\}$. We refer to the size of the implied mesh, $d_{u}$, as the dimensionality of u. We observe that $A$ is diagonalized in $U$ with eigenvalues $\{\lambda_{k}\}_{k\in\mathbb{Z}^{2}\setminus{\{0\}}}$, where $\lambda_{k}=|k|^{2}$. Periodic boundary conditions allow the use of spectral Galerkin methods when solving the Navier-Stokes PDE numerically with Fast Fourier Transforms (FFTs). When facing different boundary conditions or forcing, we would use Finite Elements techniques or other methods [21]. Convolutions arising from products in the nonlinear term are computed via FFTs on our mesh with additional antialiasing using double-sized zero padding. In addition, exponential time differencing is used as in [12] to deal with the resulting stiff system, whereby an analytical integration is used for the linear part, together with an explicit numerical Euler integration scheme for the nonlinear part.
@@
Figure 2: (Stationary regime) Heat map of the vorticity (top) of the initial condition $v_{0}$ with the truth (far left) and MCMC estimates: knowing the true $\alpha=2.2$ value (centre left), underestimating with $\alpha=1.5$ (centre), overestimating with $\alpha=3$ (centre right) and estimating $\alpha$ (far right). The crosses and circles indicate the positions where the vector field is observed. Heat map of the variance of each estimate (bottom). Note the scale of the variance being inversely proportional to $\alpha$. The algorithms knowing the true $\alpha$ and estimating it via MwG exhibit similar behaviour.
The diagnostics plots for the different MwGs are in the Appendix in section C.1. PCN and MwG run times are shown in Table C. We remark that as expected the runtimes are close to that of Algorithm 1, with $\alpha,\beta^{2}$ known. In Fig. 1 we plot the trace plots and densities of our parameter estimation and the results prove the success of the parameter estimation even with uninformative priors. We now study the algorithm’s uncertainty quantification of the initial condition. In Fig. 2 we plot the vorticity of the initial condition’s estimates against the true value. We observe similar estimates from the algorithms knowing the correct $\alpha$ and estimating it. The estimates recover the true initial condition relatively accurately, albeit giving a smoother estimate.
We now focus on the variance of the estimates’ vorticity to fully evaluate the algorithm’s uncertainty quantification. We see from the scales that the variance is highly impacted by the choice of $\alpha$. We recover the order of variance when the true $\alpha$ is known only by estimating $\alpha$. When $\alpha$ is over and underestimated the uncertainty is under and over estimated, respectfully. Practitioners often choose to sacrifice uncertainty quantification for improved point estimates and vice-versa. This could be used within our framework by adjusting the priors of $\alpha$. Note however we will use a high viscosity and this lowers the system’s energy, and in doing so, this stationary regime renders our variance estimation less crucial and the predictive power of all the algorithms similar after a few timesteps. We repeat this experiment below for the ”chaotic” regime with lower viscosity and more frequent and informative observations to investigate the differences.
@@
Figure 3: (Chaotic regime) Heat map of the variance of the velocity squared of $v_{0}$ vs $\alpha$ (top), as previously arranged, with: $\alpha=2.2$ value (far left), $\alpha=1.5$ (centre left), $\alpha=3$ (centre right) and estimating $\alpha$ (far right). The white circles in the top panels indicate the positions where the vector field is observed. We also plot the trajectory of the velocity squared of those estimates at observed (middle) and unobserved (bottom) locations. The vertical dashed line indicates the time of the last observation. Note the scale of the variance of the prediction estimates being inversely proportional to $\alpha$.
@@
We obtain similarly good results in parameter estimation as in the stationary case, see Fig. 8. This is also the case for the vorticity point estimates shown in Fig. 10. Fig. 9 illustrates that when $\alpha$ is misspecified, the estimate of $\beta^{2}$ will compensate in order to recover the correct variance, leading to poor parameter estimation. These plots can be found in the Appendix in section C.2. In Fig. 3 we present the results similarly to Fig. 2 previously. The scale of the variance of the vorticity estimates is as before at unobserved locations, however there is spatial correlation introduced by the observations. The likelihood is peaky around the informative observation locations, which reduces the variance of the posterior of the velocity at these locations, as well as the posterior’s dependence on the prior. For higher values of $\alpha$ this spatial correlation is more pronounced as the corresponding prior variance is smoothed. So a higher $\alpha$ leads to a less informative prior variance, additionally increasing the posterior’s dependence on the likelihood. This is why the variance is impacted by $\alpha$ as it was in the stationary setting only at unobserved locations. This can be additionally seen in the plots of the trajectories of the velocity’s estimates at observed times. Due to the chaotic regime having high energy, the trajectory plots show that the difference in variances does not vanish in time and in turn impacts the estimate’s predictive power. $\alpha=3$ leads to an underestimated variance at unobserved locations: the confidence interval does not capture the true value after very few time steps past the observations. At observed locations the forecast performs well, meaning the likelihood is enough to correct the misspecified prior. As for $\alpha=1.5$, the variance is overestimated leading to poor uncertainty quantification at all times. At observed locations its forecast performs well until the last observation, its predictive power degrades quickly as the variance explodes. Then MwG with $\alpha=1.5$ performs poorly at uncertainty quantification, particularly for unobserved states and prediction. This shows that even where the likelihood is most influential, the prior still influences the algorithm’s predictive power. Finally the MwG estimating $\alpha$ reproduces similar performance to the algorithm assuming its correct value and performs similarly well even in prediction.
@@ C.1 Diagnostics
Figure 7: MwG estimating $\alpha$ diagnostics: the cumulative average of the acceptance ratio of the initial condition and of $\alpha$ is around $0.23$ as desired. The acceptance ratio is diverse as desired. The trace plot of the real part of the velocity in the first dimension at various points shows each MCMC estimate (blue) captures the true value (green) in its distribution, with its cumulative average (red) converging to a value mostly close to the truth.
Fig. 6 diagnostics for pCN show that $N=8\cdot 10^{5}$ iterations are sufficient for the MCMC to provide good estimates. In Fig. 6(a) we can observe that the acceptance ratio is stable after the burn-in period of $8\cdot 10^{4}$ and between the desired values of 0.2 and 0.3. Focusing on the real value of a single Fourier coefficient, in Fig. 6(b) the MCMC estimate performs well. Additionally its autocorrelation plot in Fig. 6(c) suggests that despite high decorrelation lags, the high number of MCMC iterations will lead to a reasonable effective sample size. For MwG, we increase the number of MCMC moves to the traditional MCMC algorithm to take into account the Gibbs sampler’s moves on the hyperparameters: $N=1.6\cdot 10^{6}$ for $p_{v_{0}}=p_{\beta^{2}}=\frac{1}{3}$ and $N=2.4\cdot 10^{6}$ for $p_{v_{0}}=p_{\beta^{2}}=p_{\alpha}=\frac{1}{3}$. To fully observe the behaviour of the parameter estimation, we use longer runs: $N=6\cdot 10^{6}$. This is solely for our study and would not be needed in practice. We found the optimal step-size for $\alpha$ to be $0.96$ for both regimes. Trace plots of the velocity estimates are shown in Fig. 7.
C.2 Chaotic regime plots
Figure 9: $\beta^{2}$ estimates: Trace (top) and density (bottom) plots of prior parameter $\beta^{2}$ for MwG: knowing the true $\alpha=2.2$ value (centre), underestimating with $\alpha=1.5$ (left), overestimating with $\alpha=3$ (right). The trace plots show the estimates (blue), with its cumulative average (red) and the true value (green). The density plots display the posterior’s estimate (blue) peaking around the true value (red) of both parameters. The prior corresponds to the black dashed line. The true value lies in the support of the density of the estimates for the correct $\alpha$. However this density when $\alpha=1.5$ underestimates $\beta^{2}$ and conversely for $\alpha=3$.
Figure 10: Vorticity of the initial condition $v_{0}$ with the truth (far left) and MCMC estimates: knowing the true $\alpha=2.2$ value (centre left), underestimating $\alpha=1.5$ (centre), overestimating $\alpha=3$ (centre right) and estimating $\alpha$ (far right). The crosses indicate the positions where the vector field is observed. We observe similar estimates from the algorithms knowing the correct smoothness and estimating it, albeit giving a smooth estimate.
@@ C.2 Chaotic regime plots
MwG1 Stationary N-S $N=1.6\cdot 10^{6}$ 27.5 hours $2\cdot 10^{8}$
MwG2 Stationary N-S $N=6\cdot 10^{6}$ 85 hours $2\cdot 10^{8}$
@@
MwG1 Chaotic N-S $N=1.6\cdot 10^{6}$ 15 hours $4\cdot 10^{6}$
MwG2 Chaotic N-S $N=6\cdot 10^{6}$ 49 hours $5\cdot 10^{6}$
@@
Table 1: Run times of the MwG algorithm estimating $\theta,\phi$ and of the pCN algorithm knowing $\theta,\phi$. In Case study 1, the MwG with $p_{v_{0}}=p_{\beta^{2}}=1/2$ and with $p_{v_{0}}=p_{\beta^{2}}=p_{\alpha}=1/3$, denoted respectively as MwG1 and MwG2, are slightly slower than pCN due to additional memory requirements for storing hyperparameter samples. Taking this into account and looking at the number of times the PDE solver was ran, the run times highlight that estimating $\theta$ does not significantly increase the run times over the known $\theta$ case. The computational run times are dominated by the runs of the PDE solver when evaluating the likelihood, which was confirmed when profiling the code.